# DiCo: Revitalizing ConvNets for Scalable and Efficient Diffusion Modeling

### v1

@@ 5 Conclusion
We propose a new backbone for diffusion models, Diffusion ConvNet (DiCo), as a compelling alternative to the Diffusion Transformer (DiT). DiCo replaces self-attention with a combination of pointwise convolutions and depthwise convolutions, and incorporates a compact channel attention mechanism to reduce channel redundancy and enhance feature diversity. As a fully convolutional network, DiCo surpasses state-of-the-art diffusion models on the ImageNet 256256 and 512512 benchmarks, while achieving significant efficiency gains. We look forward to further scaling up DiCo and extending it to broader generative tasks, such as text-to-image generation.
@@
Note that the generated samples in Sec. E have been compressed to avoid exceeding the paper size limit. We include the uncompressed generated samples in the submitted Supplementary Material.
@@ Appendix C Scalability Analysis
Impact of scaling on metrics. Table. 7 and Fig. 8 illustrate the impact of DiCo model scaling across various evaluation metrics. Our results show that scaling DiCo consistently enhances performance across all metrics throughout training, highlighting its potential as a strong candidate for a large-scale foundational diffusion model.
@@ Appendix D Additional Comparison Results
In this section, we report results from state-of-the-art generative models spanning various categories, including GAN-based, masked prediction-based, autoregressive, visual autoregressive, and diffusion-based approaches. It is important to emphasize that the goal of this comparison is not to surpass all existing models, but to demonstrate that our diffusion-based DiCo model achieves competitive performance within the broader generative modeling landscape.
ImageNet 256256. Table 8 presents a comparison across generative model family on ImageNet 256256. Among diffusion models, our DiCo-XL achieves a strong FID of 2.05 with only 87.3 Gflops. Our largest variant, DiCo-H, attains a state-of-the-art FID of 1.90. When compared to other generative model types, DiCo also demonstrates competitive performance. Notably, DiCo-H, with just 1B parameters, outperforms VAR-d30—which has 2B parameters—in terms of FID.
ImageNet 512512. Table 9 presents the results on ImageNet 512512. Among diffusion models, our DiCo-XL achieves a state-of-the-art FID of 2.53 with only 349.8 GFLOPs. Compared to other generative models, DiCo continues to demonstrate strong performance. Specifically, DiCo-XL, with only 701M parameters, outperforms VAR-d36-s, which has 2.3B parameters, achieving superior FID performance with significantly fewer parameters.
@@ Appendix F Limitations
While this work demonstrates the strong performance and scalability of DiCo through extensive experiments, there are some limitations to note. Due to limited computational resources, our experiments focus exclusively on class-conditional image generation on ImageNet and do not explore text-to-image generation tasks. Moreover, our model is scaled to 1B parameters, while some advanced generative models have been scaled to even larger sizes. We aim to investigate the broader generative potential of DiCo in future research.

### v2

@@ 5 Conclusion
We propose a new backbone for diffusion models, Diffusion ConvNet (DiCo), as a compelling alternative to the Diffusion Transformer (DiT). DiCo replaces self-attention with a combination of pointwise convolutions and depthwise convolutions, and incorporates a compact channel attention mechanism to reduce channel redundancy and enhance feature diversity. As a fully convolutional network, DiCo surpasses recent diffusion models on the ImageNet 256256 and 512512 benchmarks, while achieving significant efficiency gains. Furthermore, the purely convolutional DiCo demonstrates strong potential in text-to-image generation. We look forward to further scaling up DiCo and extending it to broader generative tasks.
Acknowledgements
This research is partially funded by National Natural Science Foundation of China (Grant No. 62576342), Beijing Natural Science Foundation (4252054), Youth Innovation Promotion Association CAS(Grant No.2022132), Beijing Nova Program(20230484276).
@@ Appendix B Additional Implementation Details
| Model | #Params | FLOPs | Throughput | FID | IS | Precision | Recall |
| DiT-S/2 | 33M | 241.8G | 28 it/s | 113.06 | 12.60 | 0.25 | 0.37 |
| DiCo-S | 33M | 67.8G | 143 it/s | 102.60 | 16.45 | 0.33 | 0.38 |
@@ Appendix C Scalability Analysis
Impact of scaling on metrics. Table 8 and Fig. 8 illustrate the impact of DiCo model scaling across various evaluation metrics. Our results show that scaling DiCo consistently enhances performance across all metrics throughout training, highlighting its potential as a strong candidate for a large-scale foundational diffusion model.
@@ Appendix D Additional Comparison Results
ImageNet 256256. Table 9 presents a comparison across generative model family on ImageNet 256256. Among diffusion models, our DiCo-XL achieves a strong FID of 2.05 with only 87.3 Gflops. Our largest variant, DiCo-H, attains an FID of 1.90. When compared to other generative model types, DiCo also demonstrates competitive performance. Notably, DiCo-H, with just 1B parameters, outperforms VAR-d30—which has 2B parameters—in terms of FID.
ImageNet 512512. Table 10 presents the results on ImageNet 512512. Among diffusion models, our DiCo-XL achieves an FID of 2.53 with only 349.8 GFLOPs. Compared to other generative models, DiCo continues to demonstrate strong performance. Specifically, DiCo-XL, with only 701M parameters, outperforms VAR-d36-s, which has 2.3B parameters, achieving superior FID performance with significantly fewer parameters.
ImageNet 10241024. We conduct experiments on ImageNet 10241024. Since the original DiT paper does not report results at this resolution, we train both DiT and DiCo from scratch under the same settings for a fair comparison. We use small-scale models and train them for 400K iterations. Table 11 clearly shows that DiCo scales much more efficiently to high resolutions—achieving a 5 speedup with better generation quality.
Comparison with FlowDCN. Deformable convolution is a strong and widely used convolutional variant, and FlowDCN [85] represents an important baseline in this space. To fairly compare with FlowDCN, we retain its original isotropic architecture and training setup, and replace its MultiScale DCN with our proposed Conv Module. Table 12 shows that our Conv Module achieves better generative performance and 2.1 higher throughput compared to MultiScale DCN, despite being structurally simpler and easier to implement.
Comparison with Modern DiTs. Given the rapid advancements in modern DiTs, it is important to compare our method with recent DiTs such as LightningDiT [93]. We find that after aligning with their experimental settings, our method achieves competitive performance along with a clear efficiency advantage. Table 13 shows that our model achieves an FID score of 1.33 and offers a 2.8 throughput gain over modern DiTs. This demonstrates that our convolutional backbone can match or exceed the generative quality of modern DiTs while being significantly more efficient.
@@ Appendix F Limitations
While this work demonstrates the strong performance and scalability of DiCo through extensive experiments, there are some limitations to note. Due to limited computational resources, our model is scaled to 1B parameters, while some advanced generative models have been scaled to even larger sizes. We aim to investigate the broader generative potential of DiCo in future research.